OpenStudy (anonymous):

Help solve.. 3/a+2/b=1/c

OpenStudy (saifoo.khan):

Solve for?

OpenStudy (anonymous):

A Sorry

OpenStudy (anonymous):

in terms of b and c?

OpenStudy (anonymous):

It just says solve for a..

OpenStudy (lgbasallote):

\[\large \mathbf{\frac{3}{a} + \frac{2}{b} = \frac{1}{c}}\] right? hmmm..

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok well then it shouldn't be that bad: the first thing we want to do it get anything with an "a" in it together. so we would do: 3/a= [(1/c) -(2/b)] The reason i put it in brackets is so that we traet the tire thing as one piece. then we do some cross multiplication and division: 3= [(1/c) -(2/b)]*a 3/[(1/c) -(2/b)] = a

OpenStudy (anonymous):

This also can be simplified to get \[ a=\frac{3 b c}{b-2 c} \]